OpenStudy (anonymous):

5-(a-5/a+4)=a^2-7/a+4 5-(a-5/a+4)=a^2-7/a+4 @Mathematics

OpenStudy (anonymous):

is the answer 12?

OpenStudy (agreene):

\[5-\frac{a-5}{a+4}=\frac{a^2-7}{a+4}\] is that the question?

OpenStudy (anonymous):

yup thats it

OpenStudy (agreene):

the answer isnt 12 O.o

OpenStudy (agreene):

start by multiplying both sides by a+4 and you get to \[5(a+4)-a+5=a^2-7\] \[5a+20-a+5=a^2-7\] \[4a+25=a^2-7\] \[-a^2+4a+32=0\] \[a^2-4a-32=0\] \[a^2-4a=32\] add 4 to both sides so we can factor \[a^2-4a+4=32\] \[(a-2)^2=32\] sqrt both sides \[|a-2|=6\] now we need to eliminate the absolute value: a-2=-6 or a-2=6 a=-4 or a=8 plug back in and you will see -4 doesnt work so the only solution is \[\large a=8\]

OpenStudy (anonymous):

i was way off lol

OpenStudy (agreene):

well, that one was a bit tricky--much harder than the 1st one I answered, lol

OpenStudy (anonymous):

yea i had like a whole sheet of paper with the problem worked out lol